# LiDAR, GNSS and IMU Sensor Alignment through Dynamic Time Warping to Construct 3D City Maps

### v2

@@ Abstract
LiDAR-based 3D mapping suffers from cumulative drift causing global misalignment, particularly in GNSS-constrained environments. To address this, we propose a unified framework that fuses LiDAR, GNSS, and IMU data for high-resolution city-scale mapping. The method performs velocity-based temporal alignment using Dynamic Time Warping and refines GNSS and IMU signals via extended Kalman filtering. Local maps are built using Normal Distributions Transform-based registration and pose graph optimization with loop closure detection, while global consistency is enforced using GNSS-constrained anchors followed by fine registration of overlapping segments. We also introduce a large-scale multimodal dataset captured in Perth, Western Australia to facilitate future research in this direction. Our dataset comprises 144,000 frames acquired with a 128-channel Ouster LiDAR, synchronized RTK-GNSS trajectories, and MEMS-IMU measurements across 21 urban loops. To assess geometric consistency, we evaluated our method using alignment metrics based on road centerlines and intersections to capture both global and local accuracy. The proposed framework reduces the average global alignment error from 3.32 m to 1.24 m, achieving a 61.4% improvement, and significantly decreases the intersection centroid offset from 13.22 m to 2.01 m, corresponding to an 84.8% enhancement. The constructed high-fidelity map is publicly available through IEEE Dataport and its visualization can be viewed in the provided Demo. This dataset and method together establish a new benchmark for evaluating 3D city mapping in GNSS-constrained environments.
@@ I Introduction
Light Detection and Ranging (LiDAR) sensor provides dense 3D data with rich geometric features that are robust to lighting and weather conditions, making it ideal for mapping complex urban environments [9]. LiDAR can accurately capture features such as curbs, intersections, poles, road markings, traffic signs, building facades, sidewalks, overpasses, and street furniture. These capabilities make it a key enabler of precise city-scale 3D mapping. Recent advances, including the integration of solid-state LiDAR and IMUs [10], fusion with high-precision GNSS [11, 12], and digital LiDAR architectures such as Ouster multibeam sensors [13] have contributed to the advancement of 3D city mapping in dense urban environments [14, 15, 16].
@@
Alongside the proposed framework, we present a high-resolution multisensor dataset collected in the Perth Central Business District (CBD), Western Australia. The dataset comprises 144,000 LiDAR frames from a 128-channel Ouster sensor, accompanied by synchronized RTK-GNSS and IMU measurements across 21 road loops, covering approximately the $4.2,\text{km}^{2}$ region. LiDAR data are preprocessed through Statistical Outlier Removal (SOR) for denoising and voxel grid filtering for downsampling. GNSS and IMU streams are processed with an Extended Kalman Filter and temporally aligned via DTW to ensure consistency across modalities. This dataset provides dense 3D geometry with precise georeferencing and supports benchmarking of city mapping in GNSS-denied or obstructed environments.
@@
    We propose a method for constructing 3D city maps using non-synchrous LiDAR, GNSS and IMU data.
@@
    We propose a unified sensor fusion pipeline, coined LIGMA, that integrates LiDAR, GNSS, and inertial data using a novel velocity-based alignment and multi-stage registration strategy.
@@
    We present a high-resolution dataset captured with 128 channel Ouster LiDAR covering $4.2\,\text{km}^{2}$ of central Perth in Western Australia. The dataset comprises 144,000 LiDAR frames along with synchronized and georeferenced inertial and positioning data.
@@ II-E Quality Assurance and Handling of Imperfections
A multi-stage quality assurance pipeline was implemented to ensure data reliability for SLAM, sensor fusion, and urban perception tasks. During collection, real-time diagnostics monitored LiDAR return rates, satellite visibility (SV count), IMU status, and RTK fix quality via onboard dashboards. Sequences with degraded signals (e.g., SV count ¡ 6 or RTK loss) were flagged for review. Data were captured over multiple low-traffic days to minimize dynamic occlusions. Post-processing included statistical and radius-based filtering to clean LiDAR noise and sparsity, while frames with excessive dropout (¿25%) were annotated for exclusion. GNSS trajectories were smoothed using an extended Kalman filter to correct jump artifacts. Temporal alignment was verified by comparing LiDAR odometry with GNSS–IMU dead reckoning, with sequences exceeding 0.5 m residual or 200 ms lag corrected or removed. Calibration was validated through rigid-body residuals and loop closure consistency; loops with drift over 0.3 m were reprocessed or discarded. All quality flags and exclusion masks are stored in YAML metadata for selective processing and robustness evaluation.
@@ III Methods
This section presents the LIGMA framework for high-precision 3D urban mapping. It leverages tightly coupled integration of LiDAR, GNSS, and IMU data. As shown in Fig. 5, the proposed method comprises four main stages: (1) multimodal preprocessing, including noise filtering, voxel downsampling, interpolation, and state estimation via an Extended Kalman Filter; (2) temporal alignment of LiDAR and IMU streams using velocity-based matching and Dynamic Time Warping; (3) hierarchical mapping through NDT registration, GNSS-constrained loop closure, and ICP refinement; and (4) quantitative evaluation using skeleton-based centerline extraction and geometric metrics such as RMSE and Hausdorff distance. The following subsections describe each stage in detail.
@@ III-A2 GNSS Data
GNSS data sampled at 5 Hz are preprocessed to ensure trajectory reliability in complex urban environments. EKF smoothing is applied to estimate the state vector, which includes position, velocity, orientation, and sensor biases [41]. The prediction step integrates the IMU data using the non-linear transition model:
$\hat{\mathbf{x}}_{t|t-1}=\hat{\mathbf{x}}_{t-1|t-1}+\int_{t-1}^{t}\mathbf{F}(\hat{\mathbf{x}}_{\tau},\mathbf{u}_{\tau})\,d\tau\vskip-2.84526pt$ (1)
while the update step incorporates GNSS observations via $\hat{\mathbf{x}}_{t|t}=\hat{\mathbf{x}}_{t|t-1}+\mathbf{K}_{t}\left(\mathbf{z}_{t}-h(\hat{\mathbf{x}}_{t|t-1})\right)$, where $\hat{\mathbf{x}}_{t|t}$ is the updated state estimate, $\mathbf{K}_{t}$ is the Kalman gain, $\mathbf{z}_{t}$ is the GNSS observation, and $h(\cdot)$ is the measurement function. To improve accuracy, outlier removal is performed using Mahalanobis distance thresholding, which excludes GNSS measurements with a distance greater than 3.0, often caused by multipath effects or weak satellite geometry [42, 43]. For brief GNSS outages (less than one second), interpolation and smoothing are applied using cubic Hermite splines to fill gaps, followed by a low-pass Butterworth filter to suppress high-frequency noise [44]. The resulting data is saved in synchronized .csv files containing cleaned position, velocity, and RTK status.
@@ III-A3 IMU Data
The IMU provides measurements of angular velocity and linear acceleration. The raw IMU data are corrected and integrated using a quaternion-based inertial navigation algorithm to estimate orientation and translational velocity. Quaternions offer numerically stable roll–pitch–yaw estimation and avoid singularities such as gimbal lock. Temporal alignment is achieved using a Pulse-Per-Second (PPS) signal from the GNSS receiver. To match the 10 Hz sampling rate of the LiDAR and the 5 hz of GNSS streams, the 73 Hz IMU data are downsampled by selecting, for each LiDAR frame, the temporally closest IMU reading. This approach minimizes skew while preserving high-frequency motion characteristics at the resolution required for sensor fusion. The calibrated motion and orientation outputs are stored in a unified .csv format for downstream multimodal processing.
@@ III-B1 LiDAR Velocity Estimation
To estimate LiDAR velocity, consecutive point cloud frames are registered using ICP algorithm, which yields a $\times4$ transformation matrix representing relative motion. The translational displacement $(\Delta x,\Delta y,\Delta z)$ extracted from this matrix is used to compute instantaneous speed as $v_{\text{LiDAR}}(t)=\frac{\sqrt{\Delta x^{2}+\Delta y^{2}+\Delta z^{2}}}{\Delta t}$, where $\Delta t$ is the time interval between frames [45].
@@ III-B2 IMU Velocity Estimation
III-B3 Temporal Alignment with DTW
Scalar velocity profiles from LiDAR and IMU are temporally aligned using Dynamic Time Warping, which computes an optimal warping path to minimize the cumulative distance between sequences. This compensates for sampling rate mismatches and time offsets, resulting in synchronized sensor streams with unified timestamps for downstream fusion and mapping.
III-B4 Temporal Alignment Methodology
Temporal alignment between LiDAR and IMU velocity profiles is achieved by detecting salient features such as prolonged stationary periods, sharp accelerations, and distinct velocity peaks. Stationary intervals, defined as segments where velocity remains below 0.1m/s for more than 2 seconds, serve as reliable temporal anchors. Dynamic Time Warping (DTW) aligns the two sequences $v_{\text{LiDAR}}(t)$ and $v_{\text{IMU}}(t)$ by minimizing cumulative squared differences:
$\begin{split}D(i,j)&=\bigl(v_{\text{LiDAR}}(i)-v_{\text{IMU}}(j)\bigr)^{2}\\ &\quad+\min\{D(i-1,j),\,D(i,j-1),\,D(i-1,j-1)\}.\end{split}$ (3)
Here, $D(i,j)$ represents the cost matrix that accumulates the minimum path cost by evaluating all valid alignment paths. This non-linear matching accounts for sampling rate differences, time offsets, and sensor noise. It results in synchronized velocity profiles with unified timestamps, enabling accurate multimodal fusion and 3D reconstruction.

### v3

# LiDAR, GNSS and IMU Sensor Fine Alignment through Dynamic Time Warping to Construct 3D City Maps

@@ Abstract
LiDAR-based 3D mapping suffers from cumulative drift causing global misalignment, particularly in GNSS-constrained environments. To address this, we propose a unified framework that fuses LiDAR, GNSS, and IMU data for high-resolution city-scale mapping. The method performs velocity-based temporal alignment using Dynamic Time Warping and refines GNSS and IMU signals via extended Kalman filtering. Local maps are built using Normal Distributions Transform-based registration and pose graph optimization with loop closure detection, while global consistency is enforced using GNSS-constrained anchors followed by fine registration of overlapping segments. We also introduce a large-scale multimodal dataset captured in Perth, Western Australia to facilitate future research in this direction. Our dataset comprises 144,000 frames acquired with a 128-channel Ouster LiDAR, synchronized RTK-GNSS trajectories, and MEMS-IMU measurements across 21 urban loops. To assess geometric consistency, we evaluated our method using alignment metrics based on road centerlines and intersections to capture both global and local accuracy. The proposed framework reduces the average global alignment error from 3.32 m to 1.24 m, achieving a 61.4% improvement, and significantly decreases the intersection centroid offset from 13.22 m to 2.01 m, corresponding to an 84.8% enhancement. The constructed high-fidelity map and raw dataset are publicly available through IEEE Dataport and its visualization can be viewed in the provided Demo. This dataset and method together establish a new benchmark for evaluating 3D city mapping in GNSS-constrained environments, with source code available at GitHub Repository.
@@ I Introduction
Light Detection and Ranging (LiDAR) sensor provides dense 3D data with rich geometric features that are robust to lighting and weather conditions, making it ideal for mapping complex urban environments [58]. LiDAR can accurately capture features such as curbs, intersections, poles, road markings, traffic signs, building facades, sidewalks, overpasses, and street furniture. These capabilities make it a key enabler of precise city-scale 3D mapping. Recent advances, including the integration of solid-state LiDAR and IMUs [42], fusion with high-precision GNSS [37, 45, 52], and digital LiDAR architectures such as Ouster multibeam sensors [30] have contributed to the advancement of 3D city mapping in dense urban environments [40, 12, 54].
@@
Two representative approaches to multimodal sensor integration are batch spatiotemporal calibration using continuous-time factor graphs and latent-space sequence alignment using Canonical Time Warping. The first approach jointly estimates a continuous-time trajectory and a single inter-sensor delay parameter from camera, LiDAR, and IMU factors [44]. These formulations assume a constant offset during a session and require sufficient six-degree-of-freedom excitation to render the time parameter observable, which is often violated in long urban drives with limited roll and pitch change. The normal equations grow with the number of knots and measurements, which limits scalability on city-scale logs. Moreover, continuous-time LiDAR odometry corrects intra-scan motion distortion but does not align independent IMU and GNSS clocks when the delay is nonstationary over time [50, 23, 7]. The second approach, including Canonical Time Warping and its deep learning variants, learns projections that map two sequences into a shared latent space and then performs alignment in that space [62, 49, 51]. These methods are effective for high-dimensional paired observations such as speech features or articulated motion, yet LiDAR–IMU synchronization provides low-dimensional kinematic cues like speed and heading, cross-view correlations degrade under long stops and sensor dropouts, and learning the projections increases memory and runtime on long urban loops when task-specific training data are unavailable.
@@
Alongside the proposed framework, we present a high-resolution multisensor dataset collected in the Perth Central Business District (CBD), Western Australia. The dataset comprises 144,000 LiDAR frames from a 128-channel Ouster sensor, accompanied by synchronized RTK-GNSS and IMU measurements across 21 road loops, covering approximately the $4.2,\text{km}^{2}$ region. LiDAR data are preprocessed through Statistical Outlier Removal (SOR) for denoising and voxel grid filtering for downsampling. GNSS and IMU streams are processed with an Extended Kalman Filter and temporally aligned via DTW to ensure consistency across modalities. This dataset provides dense 3D geometry with precise georeferencing and supports benchmarking of city mapping in GNSS-denied or obstructed environments. DTW performs well when the motion signal contains stop–go events and speed transients, but the cost surface becomes ambiguous under near-constant speed or prolonged stops. We therefore restrict warping with a Sakoe–Chiba band and seed stationary anchors; when the windowed speed variance is low we fall back to a constant-shift estimate.
@@
    We propose a method for constructing 3D city maps using non-synchrous LiDAR, GNSS and IMU data and its visualization can be viewed in the provided Demo.
@@
    We propose a unified sensor fusion pipeline, coined LIGMA, that integrates LiDAR, GNSS, and inertial data using a novel velocity-based alignment and multi-stage registration strategy. Source code has been available at GitHub Repository.
@@
    We present a high-resolution dataset captured with 128 channel Ouster LiDAR covering $4.2\,\text{km}^{2}$ of central Perth in Western Australia. The dataset comprises 144,000 LiDAR frames along with synchronized and georeferenced inertial and positioning data. The constructed high-fidelity map and raw dataset are publicly available through IEEE Dataport
@@ II-E Quality Assurance and Handling of Imperfections
A multi-stage quality assurance pipeline was implemented to ensure data reliability for SLAM, sensor fusion, and urban perception tasks. During collection, real-time diagnostics monitored LiDAR return rates, satellite visibility (SV count), IMU status, and RTK fix quality via onboard dashboards. Sequences with degraded signals (e.g., SV count <6<6 or RTK loss) were flagged for review. Data were captured over multiple low-traffic days to minimize dynamic occlusions. Post-processing included statistical and radius-based filtering to clean LiDAR noise and sparsity, while frames with excessive dropout ($25\%$) were annotated for exclusion. GNSS trajectories were smoothed using an extended Kalman filter to correct jump artifacts. Temporal alignment was verified by comparing LiDAR odometry with GNSS–IMU dead reckoning, with sequences exceeding 0.5 m residual or 200 ms lag corrected or removed. Calibration was validated through rigid-body residuals and loop closure consistency; loops with drift over 0.3 m were reprocessed or discarded. All quality flags and exclusion masks are stored in YAML metadata for selective processing and robustness evaluation.
@@ III Methods
This section presents the LIGMA framework for high-precision 3D urban mapping. It leverages tightly coupled integration of LiDAR, GNSS, and IMU data. As shown in Fig. 5, the proposed method comprises four main stages: (1) multimodal preprocessing, including noise filtering, voxel downsampling, interpolation, and state estimation via an Extended Kalman Filter; (2) temporal alignment of LiDAR and IMU streams using velocity-based matching and Dynamic Time Warping; (3) hierarchical mapping through NDT registration, GNSS-constrained loop closure, and ICP refinement; and (4) quantitative evaluation using skeleton-based centerline extraction and geometric metrics such as RMSE and Hausdorff distance. The following subsections describe each stage in detail. We adopt a loosely coupled design: GNSS contributes as soft factors in the pose graph, while temporal alignment is handled in preprocessing. This keeps the synchronizer independent of GNSS availability in urban canyons and simplifies deployment at scale. Tightly coupled LiDAR–IMU–GNSS back-ends can yield further improvements when raw measurements are accessible, but they require different interfaces and formulations [46, 8].
@@ III-A2 GNSS Data
GNSS data sampled at 5 Hz are preprocessed to ensure trajectory reliability in complex urban environments. As shown in Fig. 5, we employ an Extended Kalman Filter (EKF) as a preprocessing step to (i) smooth the GNSS trajectory under multipath and intermittent RTK degradation and (ii) provide a bias-aware inertial dead-reckoning estimate during short GNSS gaps. The EKF is implemented in a local ENU navigation frame and is propagated at the IMU rate (73 Hz) with GNSS updates applied at 5 Hz. EKF smoothing is applied to estimate the state vector, which includes position, velocity, orientation, and sensor biases [47].
Specifically, the EKF state at epoch $k$ is
$\mathbf{x}_{k}=\Big[\mathbf{p}_{k}^{\top},\ \mathbf{v}_{k}^{\top},\ \mathbf{q}_{k}^{\top},\ \mathbf{b}_{g,k}^{\top},\ \mathbf{b}_{a,k}^{\top}\Big]^{\top},$ (1)
where $\mathbf{p}_{k}\in\mathbb{R}^{3}$ and $\mathbf{v}_{k}\in\mathbb{R}^{3}$ are position and velocity in ENU, $\mathbf{q}_{k}$ is the unit quaternion representing attitude (body-to-ENU), and $\mathbf{b}_{g,k}$ and $\mathbf{b}_{a,k}$ are gyroscope and accelerometer biases. Let $\boldsymbol{\omega}_{k}$ and $\mathbf{a}_{k}$ denote the raw IMU angular rate and acceleration at time step $\Delta t$. The discrete-time propagation follows the standard strapdown mechanization:
$\mathbf{p}_{k+1}$ $=\mathbf{p}_{k}+\mathbf{v}_{k}\Delta t+\tfrac{1}{2}\Big(\mathbf{R}(\mathbf{q}_{k})\big(\mathbf{a}_{k}-\mathbf{b}_{a,k}\big)-\mathbf{g}\Big)\Delta t^{2},$ (2)
$\mathbf{v}_{k+1}$ $=\mathbf{v}_{k}+\Big(\mathbf{R}(\mathbf{q}_{k})\big(\mathbf{a}_{k}-\mathbf{b}_{a,k}\big)-\mathbf{g}\Big)\Delta t,$ (3)
$\mathbf{q}_{k+1}$ $=\mathbf{q}_{k}\otimes\exp\!\Big(\tfrac{1}{2}\big(\boldsymbol{\omega}_{k}-\mathbf{b}_{g,k}\big)\Delta t\Big),$ (4)
$\mathbf{b}_{g,k+1}$ $=\mathbf{b}_{g,k}+\mathbf{n}_{bg,k},\qquad\mathbf{b}_{a,k+1}=\mathbf{b}_{a,k}+\mathbf{n}_{ba,k},\endlinenomath$ (5)
where $\mathbf{R}(\mathbf{q}_{k})$ converts a quaternion to a rotation matrix, $\mathbf{g}=[0,0,9.81]^{\top}$ m/s2 in ENU, $\otimes$ denotes quaternion multiplication, and $\exp(\cdot)$ maps an incremental rotation vector to a unit quaternion. The bias terms follow a random-walk model with driving noise $\mathbf{n}_{bg,k}$ and $\mathbf{n}_{ba,k}$. The covariance is propagated as
$\mathbf{P}_{k+1|k}=\mathbf{F}_{k}\,\mathbf{P}_{k|k}\,\mathbf{F}_{k}^{\top}+\mathbf{Q}_{k},$ (6)
where $\mathbf{F}_{k}$ is the linearized state transition and $\mathbf{Q}_{k}$ is constructed from the IMU noise characteristics provided by the sensor/SDK configuration (gyro/accel noise and bias random-walk).
For the measurement update, we use the GNSS position and Doppler-derived velocity reported by the uINS solution. The observation is
$\mathbf{z}_{k}=\Big[(\mathbf{p}^{\text{GNSS}}_{k})^{\top},\ (\mathbf{v}^{\text{GNSS}}_{k})^{\top}\Big]^{\top},\qquad h(\mathbf{x}_{k})=\Big[\mathbf{p}_{k}^{\top},\ \mathbf{v}_{k}^{\top}\Big]^{\top},$ (7)
with a standard EKF update $\hat{\mathbf{x}}_{k|k}=\hat{\mathbf{x}}_{k|k-1}+\mathbf{K}_{k}(\mathbf{z}_{k}-h(\hat{\mathbf{x}}_{k|k-1}))$ and $\mathbf{K}_{k}=\mathbf{P}_{k|k-1}\mathbf{H}^{\top}(\mathbf{H}\mathbf{P}_{k|k-1}\mathbf{H}^{\top}+\mathbf{R}_{k})^{-1}$, where $\mathbf{H}$ is the Jacobian of $h(\cdot)$ and $\mathbf{R}_{k}$ is taken from the GNSS-reported covariance when available.
To improve accuracy, outlier removal is performed using Mahalanobis distance thresholding, which excludes GNSS measurements with a distance greater than 3.0, often caused by multipath effects or weak satellite geometry [55, 11]. For brief GNSS outages (less than one second), interpolation and smoothing are applied using cubic Hermite splines to fill gaps, followed by a low-pass Butterworth filter to suppress high-frequency noise [13]. The resulting data is saved in synchronized .csv files containing cleaned position, velocity, and RTK status.
@@ III-A3 IMU Data
The IMU provides measurements of angular velocity and linear acceleration. The raw IMU data are corrected and integrated using a quaternion-based inertial navigation algorithm to estimate orientation and translational velocity. Quaternions offer numerically stable roll–pitch–yaw estimation and avoid singularities such as gimbal lock. Temporal alignment is achieved using a Pulse-Per-Second (PPS) signal from the GNSS receiver. To match the 10 Hz sampling rate of the LiDAR and the 5 hz of GNSS streams, the 73 Hz IMU data are downsampled by selecting, for each LiDAR frame, the temporally closest IMU reading. This approach minimizes skew while preserving high-frequency motion characteristics at the resolution required for sensor fusion. The calibrated motion and orientation outputs are stored in a unified .csv format for downstream multimodal processing. The GNSS and IMU measurements are produced by the same RUG-3 IMX-5 unit and are hardware timestamped on a single PPS clock. Their mutual alignment is therefore direct and does not require DTW. DTW is applied only between the OS1-128 LiDAR stream (10 Hz) and IMU-derived kinematics (73 Hz).
@@ III-B1 LiDAR Velocity Estimation
To estimate LiDAR velocity, consecutive point cloud frames are registered using ICP algorithm, which yields a $\times4$ transformation matrix representing relative motion. The translational displacement $(\Delta x,\Delta y,\Delta z)$ extracted from this matrix is used to compute instantaneous speed as $v_{\text{LiDAR}}(t)=\frac{\sqrt{\Delta x^{2}+\Delta y^{2}+\Delta z^{2}}}{\Delta t}$, where $\Delta t$ is the time interval between frames [18]. We deliberately avoid using RTK–GNSS 3D velocity for temporal alignment because in dense urban blocks the RTK status often toggles (FIX/FLT) and the 5 Hz velocity stream exhibits short gaps and smoothing latency, whereas LiDAR- and IMU-derived speed are continuously available at higher rates. Separately, keeping synchronization independent of GNSS (which is later used for global geo-anchoring) prevents data leakage and avoids optimistic bias. Note that RTK-GNSS 3D velocity is not used for DTW because in dense urban canyons FIX/FLT transitions, multipath and short outages introduce rate and latency inconsistencies [15, 53]. GNSS is retained only as a global geo-anchor. Yaw is informative but enters the DTW cost as a low-weight auxiliary term, while scalar speed remains the primary alignment cue [16].
@@ III-B2 IMU Velocity Estimation
III-B3 Linear Time-Shift Baseline
We estimate a time linear offset $\hat{\tau}$ by maximizing the normalized cross-correlation between speed profiles $v_{\text{LiDAR}}(t)$ and $v_{\text{IMU}}(t)$:
$\hat{\tau}=\arg\max_{\tau}\frac{\sum_{t}\bigl(v_{\text{LiDAR}}(t)-\bar{v}_{L}\bigr)\,\bigl(v_{\text{IMU}}(t+\tau)-\bar{v}_{I}\bigr)}{\sigma_{L}\,\sigma_{I}}$ (9)
[…]
III-B4 DTW-based Synchronization
Let $v_{L}[i]$ and $v_{I}[j]$ denote LiDAR- and IMU-derived speed sequences resampled at 10 Hz, and let $\psi_{L}[i]$ and $\psi_{I}[j]$ be the corresponding headings (IMU yaw projected to the LiDAR frame). We align them using Dynamic Time Warping (DTW) with a Sakoe–Chiba band $|i-j|\leq W$ (we use $W=25$, i.e., 2.5 s at 10 Hz) and a speed+heading cost:
$c(i,j)=\bigl(v_{L}[i]-v_{I}[j]\bigr)^{2}+\lambda\,\operatorname{wrap}_{\pi}\!\bigl(\psi_{L}[i]-\psi_{I}[j]\bigr)^{2},$ (10)
where $\operatorname{wrap}_{\pi}(\cdot)$ maps angles to $(-\pi,\pi]$ and we set $\lambda=0.25$. The cumulative cost obeys
$D(i,j)=c(i,j)+\min\{D(i-1,j),\,D(i,j-1),\,D(i-1,j-1)\},$ (11)
with monotonicity and continuity constraints; the complexity is $\mathcal{O}(NW)$. To improve stability, we seed anchors by forcing matches on stationary segments where both speeds are below $0.1\,\mathrm{m/s}$ for at least $2$ s. The optimal warping path $\mathcal{W}=\{(i_{k},j_{k})\}_{k=1}^{K}$ yields a piecewise-linear mapping between timestamps $(t_{L}[i_{k}],t_{I}[j_{k}])$, from which unified timestamps are obtained via linear interpolation. The resulting time-aligned sequences are then used for GNSS/IMU anchoring and LiDAR registration downstream.